### v1

@@ 1 Introduction
How does one construct phylogenetic networks? Traditional phylogenetic inference methods can be grouped into model-based methods (e.g. Bayesian inference [WYN16, HRNB01], maximum likelihood [WYZN18, Bro01]) or non-model-based methods (e.g. distance-based [BM04], maximum parsimony [Swo98], and combinatorial [HS11]). In all cases, one must evaluate the accuracy of the output. One way to do so is to start with some correct phylogeny $N$, obtain the necessary input data from it, infer an output phylogeny $N^{\prime}$, and measure some notion of a distance between $N$ and $N^{\prime}$. Existing metrics such as rearrangement moves suffer from computational intractability [DHJ+97, BLS17, JJE+18]; others like the triplet distance [Dob75] suffer from non-identifiability (two distinct networks could be at a distance 0. See e.g., Figure 19 in [CLRV08]).
One way of avoiding these situations is to first find complete graph invariants, sometimes called encodings, for specific classes of phylogenetic networks. Statements of the sort ‘two networks are isomorphic if and only if they have the same encodings’ are typically sought after in this area. Taking the symmetric distance of the invariants often leads to a distance metric, by definition of complete graph invariants [KV08]. In this paper, we consider invariants based on so-called $\mu$-vectors. For every vertex in the graph, these vectors encode the number of paths from it to every leaf. The multiset of all $\mu$-vectors is called the $\mu$-representation of the network. Cardona et al. introduced the $\mu$-representation and showed that it can be used as a distance metric for binary tree-child networks [CRV08].
@@
Building upon the work of Cardona et al., Erdős et al. sought to extend the application of $\mu$-representations to a larger class of phylogenetic networks in [ESS19], which they dubbed orchard networks. They proposed that it was possible to determine certain subgraphs called cherries and reticulated cherries by the $\mu$-representation. Then, by a process of iteratively simplifying the network, they proposed it would be possible to find a sequence of these (reticulated) cherries from which one could uniquely reconstruct the network. It is important to note that they considered only binary networks. However, some of their findings were later refuted by Bai et al. [BESS21] who showed that it is not possible to determine reticulated cherries by the $\mu$-representation for general binary orchard networks. In that paper, Bai et al. proposed a stack-freeness constraint within the class of orchard networks to establish the encoding result. They then aimed to show encoding holds for semi-binary stack-free orchard networks (because binary networks can be made stack-free by identifying stacks of reticulations, which makes the network semi-binary), but their proof only works for networks which are binary, as Murakami showed in [Mur21], by means of a counterexample for semi-binary networks. In [CPRC23], Cardona et al. proposed an extended $\mu$-representation, which also takes into consideration the number of paths to reticulations from each node. In the paper they showed that this extended $\mu$-representation is an encoding for binary orchard networks, lifting the stack-free condition. This modification however does not show encoding for semi-binary orchard networks as originally proposed in [BESS21], as the proof is restricted to networks which are binary. The aims of this paper are to find a large subclass of semi-binary orchard networks that is encoded by $\mu$-representations and to find variants of $\mu$-representations that form an encoding for all semi-binary orchard networks.
@@
Furthermore, in Section 4 we present a fundamental equation which governs the relationship between the in-degrees of reticulations and the $\mu$-representation of a network. We prove that such an equation exists (Proposition 1), and show how this gives rise to a system of equations on the $\mu$-vector of the root and the $\mu$-vectors and in-degrees of reticulations. We furthermore show, that for reticulation-visible networks with fixed reticulation set, the system of equations generated by Proposition 1 has a unique solution (Proposition 2). Then, we define a new class of networks called strongly reticulation-visible networks, for which there is a tree-path (a path containing only tree-nodes, which may consist of just a single tree-node) to a bridge from each child of a reticulation. A bridge is an edge which disconnects the network if cut. We show that a bridge and the lowest reticulation ancestor of that bridge in any network are uniquely determined by the $\mu$-representation (Proposition 3 and Lemma 18). We then use this to show that strongly reticulation-visible networks with the same $\mu$-representations have the same modified $\mu$-representation (Theorem 3). Finally, we conclude that strongly reticulation-visible semi-binary stack-free orchard networks are encoded in the space of semi-binary stack-free networks by the $\mu$-representation (Theorem 4). This means that the cardinality of the symmetric difference of the $\mu$-representations gives a metric between these networks. Hence, it is not necessary to include indegrees in the $\mu$-representation for this class.
@@ 2 Preliminaries
By a rooted directed acyclic graph (rooted DAG) we mean a connected directed graph that has no directed cycles. A rooted DAG whose leaves are bijectively labeled by the elements of a finite set $X$, we call an $X$-DAG. The in-degree of a node $v$ in an $X$-DAG, which we denote $\delta^{-}(v)$, is the number of edges which end in $v$. The out-degree $\delta^{+}(v)$ of a node $v$ is the number of edges starting in $v$. The degree of a node is the sum of its in-degree and out-degree. Every vertex of an $X$-DAG can be classified by their in-degree or their out-degree. In particular, an $X$-DAG contains
@@
A phylogenetic network is an $X$-DAG without parallel arcs or elementary nodes, where the root must be a leaf (in which case the network is one on a single leaf) or have out-degree greater than or equal to 2, reticulations have out-degree 1 and leaves are tree-nodes. A phylogenetic network in which all nodes except the root or the leaves have degree 3 is called binary. A phylogenetic network in which all tree-nodes except the root or the leaves have degree 3 but reticulations can have degree $\geq 3$ is called semi-binary. We say a phylogenetic network is non-binary when there are no such added restrictions on the degrees of the nodes. Note that non-binary does not mean that the network is ‘not binary’, but rather, that the network is ‘not necessarily binary’. A phylogenetic network which does not contain any reticulations is called a phylogenetic tree. See Figure 3 for some examples. Henceforth, we may refer to phylogenetic networks as networks for brevity.
@@
A path $v_{0}\rightsquigarrow v_{k}$ between two nodes $v_{0},v_{k}\in V$ is a sequence of edges $v_{0}v_{1},v_{1}v_{2},...,v_{k-1}v_{k}$ such that $v_{i}v_{i+1}\in E$ for $i\in\{0,1,\ldots,k-1\}$.
@@
A tree-path is a directed path $v_{0}\rightsquigarrow v_{k}$, such that $v_{i}$ is a tree-node for each $i\in\{0,1,\ldots,k\}$. A tree-node $v$ which has out-degree $\delta^{+}(v)=1$ we shall call an elementary node. A path for which all but the start and end nodes are elementary nodes, we shall call an elementary path. The height of a node is the length of the longest path from the node to a leaf.
@@ 2.1 Network classes
Figure 3: A binary phylogenetic tree, a binary phylogenetic network and a semi-binary phylogenetic network
@@
A network is called orchard if there exists a sequence $s_{1}s_{2}s_{3}\ldots s_{i}\dots s_{n}$, of ordered pairs, such that $s_{i}$ is a reducible pair in the network after reducing each pair in the sequence up to $s_{i-1}$ and the entire sequence reduces the network to a network on a single leaf. Note that in that case, each network generated by performing reductions $s_{1}$ up to $s_{i}$, is orchard with sequence $s_{i+1},s_{i+2},\ldots,s_{n}$, see Corollary 4.2 in [ESS19]. The network in Figure 4 is orchard with sequence $(b,c)(a,c)(b,a)(a,c)(c,a)$. It contains the reticulated cherry $(b,c)$. The networks in Figure 4 are all orchard and contain the reducible pair $(b,a)$. In the phylogenetic tree $(b,a)$ is a cherry, while in the other networks $(b,a)$ is a reticulated cherry. See Figure 5 for a visualization of the way the different classes of phylogenetic networks discussed in this paper are related.
@@
Figure 5: A venn diagram showing the relations between several different classes of phylogenetic networks
@@ 2.2 The $\mu$-representation
Given an $X$-DAG $\mathcal{N}=(V,E)$, the path-multiplicity vector or $\mu$-vector of any node $v\in V$ is defined as follows: let $\mu(v)\in\mathbb{Z}^{X}$ be a vector indexed by the leaves of $\mathcal{N}$, such that the element indexed by leaf $a$, denoted $\mu_{a}(v)$, is equal to the number of paths from $v$ to $a$. Note that $\mu(v)$ only contains non-negative integer elements and is never equal to the zero vector. Moreover, as for each leaf there is only the trivial path from that leaf to itself, the $\mu$-vectors of leaves are unit-vectors. With the exception of leaf nodes, the $\mu$-vector of a node is always the sum of the $\mu$-vectors of its children. For two vertices $u,v\in V$, we write $\mu(u)\leq\mu(v)$ if $\mu_{a}(u)\leq\mu_{a}(v)$ for all $a\in X$. We write $\mu(u)<\mu(v)$ if $\mu(u)\leq\mu(v)$ and $\mu_{a}(u)<\mu_{a}(v)$ for at least one $a\in X$. Note that if $u$ is descendant of $v$, it always holds that $\mu(u)\leq\mu(v)$. We say that a $\mu$-vector $\mu(v)$ belongs to a node $u$ if $\mu(u)=\mu(v)$. In particular, $\mu(v)$ belongs to $v$. Later on, we shall see that a $\mu$-vector can belong to multiple nodes. The $\mu$-representation of $\mathcal{N}$, denoted $\mu(\mathcal{N})$, is the multiset of all $\mu$-vectors of nodes in $V$.
@@
The multisets considered here will always be taken in the set of $\mu$-vectors. If a given $\mu$-vector $\mu(v)$ is not contained in a multiset $\mu(\mathcal{N})$, we may say $\mu(v)$ has multiplicity 0 in $\mu(\mathcal{N})$. The operation of removing a $\mu$-vector $\mu(v)$ from $\mu(\mathcal{N})$ is equivalent to lowering the multiplicity of $\mu(v)$ in $\mu(\mathcal{N})$ by 1. Clearly, the multiplicity of a $\mu$-vector cannot be negative and a $\mu$-vector which has multiplicity 0 in $\mu(\mathcal{N})$ cannot be removed from $\mu(\mathcal{N})$. The operation of adding a $\mu$-vector $\mu(v)$ to a multiset $\mu(\mathcal{N})$ is equivalent to increasing the multiplicity of $\mu(v)$ in $\mu(\mathcal{N})$ by 1. We say that $\mu(\mathcal{N})$ is generated by adding $\mu(v)$ for each node $v\in V$. Therefore, the multiplicity of a vector $\mu(v)$ in $\mu(\mathcal{N})$ is equal to the number of nodes in $\mathcal{N}$ with $\mu$-vector equal to $\mu(v)$. We do not equate the nodes $v\in V$ with their $\mu$-vectors because multiple nodes may have the same $\mu$-vector. A set of at least two tree-nodes which have the same $\mu$-vector we shall call tree-clones. A node which is part of a set of tree-clones, we shall call a tree-clone. Bai et al. showed in [BESS21] Lemma 4.4, that orchard networks do not contain tree-clones.
@@ 2.3 The symmetric difference of multisets
The cardinality of the symmetric difference, or the number of elements that are unique to either set, can be used as a measure for the difference between these two sets. For multisets the symmetric difference is defined somewhat differently.
Let $M_{1}$ and $M_{2}$ be multisets. Let $x\in M_{1}$ with multiplicity $i$ and $x\in M_{2}$ with multiplicity $j$. Then $x\in M_{1}\triangle M_{2}$ with multiplicity $|i-j|$. In other words, the symmetric difference $M_{1}\triangle M_{2}$ contains elements in $M_{1}$ or $M_{2}$, with multiplicity equal to the absolute difference of the respective multiplicities within $M_{1}$ or $M_{2}$. The cardinality of a multiset is the sum of the multiplicities of all its elements. Let $M_{3}$ be another multiset. The following equalities and inequalities hold for the symmetric difference cardinalities.
  • $|M_{1}\triangle M_{2}|\geq 0$,
  • $|M_{1}\triangle M_{2}|=0$ if, and only if $M_{1}=M_{2}$,
  • $|M_{1}\triangle M_{2}|=|M_{2}\triangle M_{1}|$, and
  • $|M_{1}\triangle M_{3}|\leq|M_{1}\triangle M_{2}|+|M_{2}\triangle M_{3}|$.
The second condition is true, because a multiset can be uniquely represented by its multiplicity function, therefore if two multisets have equal multiplicity functions then they are the same multiset. This makes the cardinality of the symmetric difference a distance function or metric on multisets. For it to be a metric on phylogenetic networks however, we need a modified version of the second condition to hold. Let $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ be networks.
  • $|\mu(\mathcal{N}_{1})\triangle\mu(\mathcal{N}_{2})|=0$ if, and only if $\mathcal{N}_{1}\cong\mathcal{N}_{2}$.
Therefore, we will be focusing on determining the conditions on $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ or on $\mu$ such that the following holds:
  • $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$ if, and only if $\mathcal{N}_{1}\cong\mathcal{N}_{2}$.
It is important to note that there is always only one $\mu$-representation belonging to a given network.

### v2

@@ 1 Introduction
How does one construct phylogenetic networks? Traditional phylogenetic inference methods can be grouped into model-based methods (e.g. Bayesian inference [WYN16, HRNB01], maximum likelihood [WYZN18, Bro01]) or non-model-based methods (e.g. distance-based [BM04], maximum parsimony [Swo98], and combinatorial [HS11]). In all cases, one must evaluate the accuracy of the output. For certain evolutionary histories, the true phylogeny is known. In validating the inference method, this means one can compare output phylogenies to the benchmark phylogeny. In doing so, one needs a notion of computing distances between the two phylogenies. Existing metrics such as rearrangement moves suffer from computational intractability [DHJ+97, BLS17, JJE+18]; others like the triplet distance [Dob75] suffer from non-identifiability (two distinct networks could be at a distance 0. See e.g., Figure 19 in [CLRV08]).
One way of avoiding these situations is to first find complete graph invariants, sometimes called encodings, for specific classes of phylogenetic networks. Statements of the sort ‘two networks are isomorphic if and only if they have the same encodings’ are typically sought after in this area. Taking the symmetric difference of the invariants often leads to a metric, by definition of complete graph invariants [Gur01]. In this paper, we consider invariants based on so-called $\mu$-vectors. For every vertex in the graph, these vectors encode the number of paths from it to every leaf. The multiset of all $\mu$-vectors is called the $\mu$-representation of the network. Cardona et al. introduced the $\mu$-representation and showed that it can be used as a metric for binary tree-child networks [CRV08].
@@
Building upon the work of Cardona et al., Erdős et al. sought to extend the application of $\mu$-representations to a larger class of phylogenetic networks in [ESS19], which they called orchard networks. Their proof was based on identifying and reducing so-called cherries and reticulated cherries, straight from the $\mu$-representations. However, some of their findings were later refuted by Bai et al. [BESS21] who showed that it is not possible to determine reticulated cherries from the $\mu$-representation for general binary orchard networks. In that paper, Bai et al. proposed a stack-freeness constraint within the class of orchard networks to establish the encoding result. They then aimed to show encoding holds for semi-binary stack-free orchard networks (because binary networks can be made stack-free by identifying stacks of reticulations, which makes the network semi-binary). This claim was shown, by counter-example, to only hold for networks which are binary [Mur21]. In [CPRC23], Cardona et al. proposed an extended $\mu$-representation, which also takes into consideration the number of paths to reticulations from each node. In the paper they showed that this extended $\mu$-representation is an encoding for binary orchard networks, lifting the stack-free condition. This modification however does not show encoding for semi-binary orchard networks as originally proposed in [BESS21], as the proof is restricted to networks which are binary.
As it stands, $\mu$-representations encode stack-free binary orchard networks, and the extended $\mu$-representations encode binary orchard networks. The aims of this paper are to find a large subclass of semi-binary orchard networks that is encoded by $\mu$-representations and to find variants of $\mu$-representations that form an encoding for all semi-binary orchard networks.
@@
Furthermore, in Section 4 we present a fundamental equation which governs the relationship between the in-degrees of reticulations and the $\mu$-representation of a network. We prove that such an equation exists (1), and show how this gives rise to a system of equations on the $\mu$-vector of the root and the $\mu$-vectors and in-degrees of reticulations. We furthermore show, that for reticulation-visible networks with fixed reticulation set, the system of equations generated by 1 has a unique solution (2). Then, we define a new class of networks called strongly reticulation-visible networks, for which there is a tree-path (a path containing only tree-nodes, which may consist of just a single tree-node) to a bridge from each child of a reticulation. A bridge is an edge which disconnects the network if cut. We show that a bridge and the lowest reticulation ancestor of that bridge in any network are uniquely determined by the $\mu$-representation (3 and 12). We then use this to show that strongly reticulation-visible networks with the same $\mu$-representations have the same modified $\mu$-representation (Theorem 3). Finally, we conclude that strongly reticulation-visible semi-binary stack-free orchard networks are encoded in the space of semi-binary stack-free networks by the $\mu$-representation (Theorem 4). This means that the cardinality of the symmetric difference of the $\mu$-representations gives a metric between these networks. Hence, it is not necessary to include indegrees in the $\mu$-representation for this class.
@@ 2 Preliminaries
By a rooted directed acyclic graph (rooted DAG) we mean a connected directed graph that has no directed cycles. A rooted DAG whose leaves are bijectively labeled by the elements of a finite set $X$, we call an $X$-DAG. We assume henceforth that there is some ordering on the elements of $X$, i.e., that $X=\{x_{1},\ldots,x_{n}\}$. This helps with defining certain terms. The in-degree of a node $v$ in an $X$-DAG, which we denote $\delta^{-}(v)$, is the number of edges which end in $v$. The out-degree $\delta^{+}(v)$ of a node $v$ is the number of edges starting in $v$. The degree of a node is the sum of its in-degree and out-degree. Every vertex of an $X$-DAG can be classified by their in-degree or their out-degree. In particular, an $X$-DAG contains
@@
A phylogenetic network is an $X$-DAG without parallel arcs or elementary nodes, where the root must be a leaf (in which case the network is one on a single leaf) or have out-degree greater than or equal to 2, reticulations have out-degree 1 and leaves are tree-nodes. A phylogenetic network in which all nodes except the root or the leaves have degree 3 is called binary. A phylogenetic network in which all tree-nodes except the root or the leaves have degree 3 but reticulations can have degree $\geq 3$ is called semi-binary. Biologically, such evolutionary histories can occur if there are ambiguities in the order of consecutive reticulate events. We say a phylogenetic network is non-binary when there are no such added restrictions on the degrees of the nodes. Note that non-binary does not mean that the network is ‘not binary’, but rather, that the network is ‘not necessarily binary’. A phylogenetic network which does not contain any reticulations is called a phylogenetic tree. See Figure 3 for some examples. Henceforth, we may refer to phylogenetic networks as networks for brevity.
@@
A path $v_{0}\rightsquigarrow v_{k}$ between two nodes $v_{0},v_{k}\in V$ is a sequence of edges $v_{0}v_{1},v_{1}v_{2},...,v_{k-1}v_{k}$ such that $v_{i}v_{i+1}\in E$ for $i\in\{0,1,\ldots,k-1\}$. Note, that this means that in this paper all paths will be directed paths as all edges are directed.
@@
A tree-path is a path $v_{0}\rightsquigarrow v_{k}$, such that $v_{i}$ is a tree-node for each $i\in\{0,1,\ldots,k\}$. A tree-node $v$ which has out-degree $\delta^{+}(v)=1$ we shall call an elementary node. A path for which all but the start and end nodes are elementary nodes, we shall call an elementary path. The height of a node is the length of the longest path from the node to a leaf.
@@ 2.1 Network classes
Figure 3: A binary phylogenetic tree, a binary phylogenetic network and a semi-binary phylogenetic network.
@@
A network is called orchard if there exists a sequence $s_{1}s_{2}s_{3}\ldots s_{i}\dots s_{n}$, of ordered pairs, such that $s_{i}$ is a reducible pair in the network after reducing each pair in the sequence up to $s_{i-1}$ and the entire sequence reduces the network to a network on a single leaf. Note that in that case, each network generated by performing reductions $s_{1}$ up to $s_{i}$, is orchard with sequence $s_{i+1},s_{i+2},\ldots,s_{n}$, see Corollary 4.2 in [ESS19]. Furthermore, any maximal sequence of cherry reductions is complete, meaning any cherry reduction will result in an orchard network and any partial sequence can be made complete. The network in Figure 4 is orchard with sequence $(b,c)(a,c)(b,a)(a,c)(c,a)$. It contains the reticulated cherry $(b,c)$. The networks in Figure 4 are all orchard and contain the reducible pair $(b,a)$. In the phylogenetic tree $(b,a)$ is a cherry, while in the other networks $(b,a)$ is a reticulated cherry. See Figure 5 for a visualization of the way the different classes of phylogenetic networks discussed in this paper are related.
@@
Figure 5: A diagram showing the relations between several different classes of phylogenetic networks.
@@ 2.2 The $\mu$-representation
Given an $X$-DAG $\mathcal{N}=(V,E)$, the path-multiplicity vector or $\mu$-vector of any node $v\in V$ is defined as follows: let $\mu(v)\in\mathbb{Z}^{X}$ be a vector indexed by the leaves of $\mathcal{N}$, such that the coordinate indexed by leaf $a$, denoted $\mu(v)_{a}$, is equal to the number of paths from $v$ to $a$. Note that $\mu(v)$ only contains non-negative integer coordinates and is never equal to the zero vector. Moreover, since any leaf constitutes a trivial path as well their $\mu$-vector is the corresponding unit vector. With the exception of leaf nodes, the $\mu$-vector of a node is always the sum of the $\mu$-vectors of its children. For two vertices $u,v\in V$, we write $\mu(u)\leq\mu(v)$ if $\mu(u)_{a}\leq\mu(v)_{a}$ for all $a\in X$. We write $\mu(u)<\mu(v)$ if $\mu(u)\leq\mu(v)$ and $\mu(u)_{a}<\mu(v)_{a}$ for at least one $a\in X$. Note that if $u$ is descendant of $v$, it always holds that $\mu(u)\leq\mu(v)$. We say that a $\mu$-vector $\mu(v)$ belongs to a node $u$ if $\mu(u)=\mu(v)$. In particular, $\mu(v)$ belongs to $v$. Later on, we shall see that a $\mu$-vector can belong to multiple nodes. The $\mu$-representation of $\mathcal{N}$, denoted $\mu(\mathcal{N})$, is the multiset of all $\mu$-vectors of nodes in $V$.
@@
If a given $\mu$-vector $\mu(v)$ is not contained in a multiset $\mu(\mathcal{N})$, we may say $\mu(v)$ has multiplicity 0 in $\mu(\mathcal{N})$. The operation of removing a $\mu$-vector $\mu(v)$ from $\mu(\mathcal{N})$ is equivalent to lowering the multiplicity of $\mu(v)$ in $\mu(\mathcal{N})$ by 1. Clearly, the multiplicity of a $\mu$-vector cannot be negative and a $\mu$-vector which has multiplicity 0 in $\mu(\mathcal{N})$ cannot be removed from $\mu(\mathcal{N})$. The operation of adding a $\mu$-vector $\mu(v)$ to a multiset $\mu(\mathcal{N})$ is equivalent to increasing the multiplicity of $\mu(v)$ in $\mu(\mathcal{N})$ by 1. We say that $\mu(\mathcal{N})$ is generated by adding $\mu(v)$ for each node $v\in V$. Therefore, the multiplicity of a vector $\mu(v)$ in $\mu(\mathcal{N})$ is equal to the number of nodes in $\mathcal{N}$ with $\mu$-vector equal to $\mu(v)$. We do not equate the nodes $v\in V$ with their $\mu$-vectors because multiple nodes may have the same $\mu$-vector. A set of at least two tree-nodes which have the same $\mu$-vector we shall call tree-clones. A node which is part of a set of tree-clones, we shall call a tree-clone. Bai et al. showed in [BESS21] Lemma 4.4, that orchard networks do not contain tree-clones.
@@ 2.3 The symmetric difference of multisets
The cardinality of the symmetric difference, or the number of elements that are unique to either set, can be used as a measure for the difference between these two sets.
For multisets $M_{1},M_{2}$, the symmetric difference $M_{1}\triangle M_{2}$ contains elements in $M_{1}$ or $M_{2}$, with multiplicity equal to the absolute difference of the respective multiplicities within $M_{1}$ or $M_{2}$. The cardinality of a multiset is the sum of the multiplicities of all its elements. The cardinality of the symmetric difference is a metric on multisets. For it to be a metric on some subset of phylogenetic networks, we need a modified version of the identity axiom to hold. Let $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ be networks.
  • $|\mu(\mathcal{N}_{1})\triangle\mu(\mathcal{N}_{2})|=0$ if, and only if $\mathcal{N}_{1}$ is isomorphic to $\mathcal{N}_{2}$.
Therefore, we will be focusing on determining the conditions such that the $\mu$-representations are equal if, and only if the networks are isomorphic. It is important to note that there is always only one $\mu$-representation belonging to a given network.